OpenStudy (anonymous):

did i simplify this right??? cos^2(theta)(1+tan^2(theta)) simplified 1 or (cos^2(theta)) (1/cos^2(theta))

OpenStudy (raden):

yes, u r right... the answer is 1

OpenStudy (anonymous):

thanks you :)

OpenStudy (raden):

yw

OpenStudy (anonymous):

can you help me with another problem???

OpenStudy (raden):

sure

OpenStudy (anonymous):

okay the given function is : \[\cot(x)\div(\csc(-x))\] i have it simplified down to\[\cos(x)(-\sin(x))divsin(x)\] and now i am stuck

OpenStudy (raden):

ok, convert cot(x) to cos(x)/sin(x) and the identity of csc(-x) = -csc(x) and dont forget csc(x) = 1/sin(x) so, cot(x) : csc(-x) = cot(x) : (-csc(x) = cos(x)/sin(x) : 1/(-sin(x) = cos(x)/sin(x) * -sin(x)/1 = -cos(x)

OpenStudy (anonymous):

ty

OpenStudy (raden):

welcome

OpenStudy (anonymous):

i wasnt sure if you could do that since the one sin was negative and the other was not

OpenStudy (raden):

nopes...

OpenStudy (anonymous):

nope as in my thinking is wronog???

OpenStudy (raden):

opsss. sory i was misscomunication :)

OpenStudy (anonymous):

so is -cos(x) the answer?

OpenStudy (raden):

yes it is

OpenStudy (raden):

any else

OpenStudy (anonymous):

yea (cos(x))/sec(x)+tan(x) --> cos^2(x)/sin(x) correct? -->=simplifies to

OpenStudy (raden):

is ur question

OpenStudy (anonymous):

did i simplify it correctly?

OpenStudy (raden):

cos(x)/(sec(x)+tan(x)) OR cos(x)/sec(x) + tan(x)

OpenStudy (raden):

which one :